6x - 5y = 11
7x + 5y = 2

Answer by jim_thompson5910(35256)

Add the equations to get

6x - 5y = 11
7x + 5y = 2
------------
13x+0y = 13

13x = 13
x = 13/13
x = 1

Now that you know that x = 1, use it to find y

6x - 5y = 11
6(1) - 5y = 11
6 - 5y = 11
-5y = 11-6
-5y = 5
y = 5/(-5)
y = -1


Answers: x = 1 and y = -1